Suppose Dorothy breeds lizards. In lizards, G codes for green bodies and g codes for brown bodies. In a separate assorting locus, Y codes for yellow tails and y codes for blue tails. Dorothy crosses two heterozygotes that produces 656 offspring.

How many of these 656 offspring are predicted to have both green bodies and blue tails?

Number of lizards with green bodies and blue tails= ???


Thank you!!!
Answer #1

Answer:

GgYy x GgYy---Parents

GgxGg = G_ (green bodies-3/4) & gg (brown bodies-1/4)

Yy x Yy = Y_ (yellow tail-3/4) & yy (blue tail-1/4)

The probability of both green bodies and blue tails = G_bb= ¾ * ¼ = 3/16

Number of both green bodies and blue tails = 3/16 * 656 = 123